500 ml D5W with 500 mg Aminophylline to infuse at 150 ml/hr (20gtts/ml). How many drops per minute?
Answer #1

Ans) ml/hr = Volume × Drop factor ÷ time in mins

150 = 500 × 20 ÷ Time

Time = 10000 ÷ 150 = 200 ÷ 60 = 3.33 mins